OpenStudy (praetorian.10):

percentages? i have: prophase 78 metaphase 115 anaphase 5 telophase 3 a total of 198 how do i work out what each is in percentage?

OpenStudy (praetorian.10):

like i really suck at maths ..... its probs easy but lol i got no clue

OpenStudy (praetorian.10):

yolo

OpenStudy (unklerhaukus):

how did you get a total of 198?

OpenStudy (praetorian.10):

i counted them ?

OpenStudy (unklerhaukus):

count all of them, (even count the telophase)

OpenStudy (praetorian.10):

78 + 115 + 5 + 3 = 198

OpenStudy (praetorian.10):

awww cr@p

OpenStudy (praetorian.10):

dang it

OpenStudy (praetorian.10):

one sec

OpenStudy (unklerhaukus):

78 + 115 + 5 + 3 = 198 + 3

OpenStudy (praetorian.10):

supposed to be 112

OpenStudy (praetorian.10):

for metaphase

OpenStudy (praetorian.10):

thanks for pointing that out :D

OpenStudy (unklerhaukus):

oh, ok prophase 78 metaphase 112 anaphase 5 telophase 3 total 78 + 112 + 5 + 3 = 198

OpenStudy (praetorian.10):

yep :)

OpenStudy (unklerhaukus):

to work out the percentage prophase \[78/198 = 0.3939... = 39.4\%\] the percentage metaphase \[115/198 = \]

OpenStudy (praetorian.10):

ohhhh let me try

OpenStudy (praetorian.10):

umm

OpenStudy (praetorian.10):

58.08

OpenStudy (praetorian.10):

right?

OpenStudy (unklerhaukus):

yes ! the percentage metaphase is 58.1% (to 1 decimal place)

OpenStudy (praetorian.10):

and ummm that means that 5/198 = 2.53 and 3/198 = 1.52

OpenStudy (praetorian.10):

(rounded up) do we round up?

OpenStudy (unklerhaukus):

yeah if the first digit the gets cut off in rounding is a digit 5-9 then round it up,

OpenStudy (praetorian.10):

yay!

OpenStudy (praetorian.10):

thank you so much @UnkleRhaukus

OpenStudy (praetorian.10):

so taking the time to help and explain and teach

OpenStudy (praetorian.10):



OpenStudy (unklerhaukus):

now lets check the results make sense the sum of the percentages of prophase, metaphase, anaphase, and telophase 39.4% + 58.1% + 2.5% + 1.5% = 101.5% with is close to 100%

OpenStudy (unklerhaukus):

oh wait , check that metaphase one again, 112/198 =

OpenStudy (praetorian.10):

56.57

OpenStudy (unklerhaukus):

yeah about 112/198 = 56.7% ( not 115/198 = 58.1% )

OpenStudy (praetorian.10):

thats right :D

OpenStudy (unklerhaukus):

so the sum is 39.4% + 56.7% + 2.5% + 1.5% =

OpenStudy (praetorian.10):

100.1

OpenStudy (unklerhaukus):

which is 100% plus the tiny rounding error

OpenStudy (praetorian.10):

I'm trying to fill out this table i was wondering what in the world comes under the "time / minutes' part on far right? do yo have any idea mate?

OpenStudy (praetorian.10):

i was like 0_o what a weird column to ask for ....

OpenStudy (unklerhaukus):

hmm, that is a strange column indeed. Did you time how long it took you to take the counts?, (im not really sure)

OpenStudy (praetorian.10):

not at all i was so confused when i saw it, thats okay i will send my lecturer a email

OpenStudy (praetorian.10):

thank you so much for all the help mate most appreciated

OpenStudy (praetorian.10):